l Sprint LTE 04:43 73% WebWorks Website Sec8-4PT1: Problem 2 Prev Up Next (2 pts) The sample mean and standard deviation from a random sample of 32 observations from a normal population were computed as- 31 and s 12. Calculate the t statistic of the test required to determine whether there is enough evidence to infer at the 4% significance level that the population mean is greater than 28. Test Statistic Preview Answers Submit Answers You have attempted this problem 0 times. You have unlimited attempts remaining. Email instructor Page generated at 03/22/2018 at 04:43am CDT
Answer #1

To test

Ho:\mu =28

vs H1:\mu >28

Test statistics

t=\frac{\bar{x}-\mu }{\frac{s}{\sqrt{n}}}=\frac{31-28}{\frac{12}{\sqrt{32}}}=1.414

P_{Value}=P{(T_{31}>t)}=P(T_{31}>1.414)=1-P(T_{31}<1.414)=0.08366

Rejectio Rule: If P value is less than \alpha then we reject Ho at \alpha% level of significance.

Here P value=0.08366 is greater than \alpha =0.04 hence we will accept Ho and conclude that there is no evidence to claim that population mean is greater than 28